OpenStudy (anonymous):

Find y in, -5y+3=2(4y+2) y=? step by step please

OpenStudy (anonymous):

Distribute into the parenthesis and solve. It's much simpler than it looks.

OpenStudy (anonymous):

Expand the brackets -5y+3=2*4y+2*2 -5y+3=8y+4 Then get the ys on one side and the constants on the other side -5y-8y=4-3 -13y=1 Then make y the subject y=-1/13

OpenStudy (anonymous):

OK, let's test that on the original system: \[-5y+3=2(4+2y)\] \[-5 (-1/13) + 3=2(2+4(-1/13))\] \[5/13+3=2(22/13)\] \[44/13=44/13\] The last equation is true, so you are correct!